# MetaEmbed: Scaling Multimodal Retrieval at Test-Time with Flexible Late Interaction

### v1

@@ Abstract
Universal multimodal embedding models have achieved great success in capturing semantic relevance between queries and candidates. However, current methods either condense queries and candidates into a single vector, potentially limiting the expressiveness for fine-grained information, or produce too many vectors that are prohibitively expensive for multi-vector retrieval. In this work, we introduce MetaEmbed, a new framework for multimodal retrieval that rethinks how multimodal embeddings are constructed and interacted with at scale. During training, a fixed number of learnable Meta Tokens are appended to the input sequence. At test-time, their last-layer contextualized representations serve as compact yet expressive multi-vector embeddings. Through the proposed Matryoshka Multi-Vector Retrieval training, MetaEmbed learns to organize information by granularity across multiple vectors. As a result, we enable test-time scaling in multimodal retrieval where users can balance retrieval quality against efficiency demands by selecting the number of tokens used for indexing and retrieval interactions. Extensive evaluations on the Massive Multimodal Embedding Benchmark (MMEB) and the Visual Document Retrieval Benchmark (ViDoRe) confirm that MetaEmbed achieves state-of-the-art retrieval performance while scaling robustly to models with 32B parameters.
@@ 1 Introduction
Multimodal embedding models play an essential role in image search (Gordo et al., 2016), visual question answering (Hu et al., 2018; Zheng et al., 2021) and visual document retrieval (Faysse et al., 2025), where models project heterogeneous inputs, such as images and texts, into a unified representation space. While existing methods, including CLIP (Radford et al., 2021), BLIP (Li et al., 2022) and SigLIP (Zhai et al., 2023) have demonstrated superior performance in cross-modal retrieval, their performance remains limited in scenarios where the inputs involve complex and diverse instructions. Thanks to recent advances in building embeddings through foundation vision-language models (VLMs), one could apply contrastive learning on the extracted embedding from the hidden states of the last layer of a VLM to learn meaningful multimodal representations while retaining pre-trained knowledge.
@@
Despite growing progress in multimodal embedding VLMs, the common practice of condensing the entire query and candidate into a single vector is not an optimal choice, as fine-grained details are lost between modalities (Yao et al., 2022; Thrush et al., 2022) and this process has theoretical limitations (Weller et al., 2025). In text retrieval, ColBERT (Khattab and Zaharia, 2020) introduced a multi-vector late interaction mechanism that retains multiple token-level embeddings and uses a lightweight scoring between query and document token representations. This approach preserves significantly more contextual information than single-vector methods while still allowing independent encoding of queries and documents, and has motivated a recent trend of devising multi-vector embeddings for multimodal retrieval (Faysse et al., 2025; Xu et al., 2025; Günther et al., 2025). Despite their effectiveness, multi-vector methods incur substantial efficiency costs in terms of index size, retrieval latency and feasibility. In these methods, each image is encoded into hundreds of patch embeddings, and each query text into several token embeddings, all of which must be stored and compared during retrieval. This results in large index sizes and slower retrieval processing. Moreover, multimodal-to-multimodal retrieval becomes impractical where both query and candidate sides have images, as similarity computation for each pair of query and candidate involves interactions between thousands of query tokens and thousands of candidate tokens, making both training and inference prohibitive due to computational demands.
@@
Figure 1: Different multimodal retrieval frameworks at training-time. Upper Left: Single vector retrieval method computes a score for each pair of query and candidate and uses contrastive objective to maximize the score for corresponding pairs. Upper Right: In multi-vector retrieval, scores are computed via maximum similarity across vector pairs and aggregated before applying the contrastive objective. Lower: MetaEmbed structures query and candidate vectors into hierarchical nested groups and trains coarse-to-fine multi-vector embeddings that enable scalable retrieval with flexible test-time control over embedding granularity.
@@
We first validate MetaEmbed on the Massive Multimodal Embedding Benchmark (MMEB) (Jiang et al., 2024) and Visual Document Retrieval Benchmarks (ViDoRe) v2 (Faysse et al., 2025; Macé et al., 2025), which represent a comprehensive suite of retrieval tests covering images, text and visual documents. Our experiments show that MetaEmbed achieves state-of-the-art retrieval performance across diverse scenarios. To further examine its generality and training scalability, we evaluate MetaEmbed with different VLM architectures and model sizes, including decoder-only models such as Qwen2.5-VL (Bai et al., 2025) and PaliGemma (Beyer et al., 2024), as well as cross-attention-based models such as Llama-3.2-Vision (Grattafiori et al., 2024). Notably, test-time scalability remains effective even at the 32B scale, showing minimal diminished returns with larger models. We hope MetaEmbed charts a path toward multimodal retrieval systems that are both accurate and deployable at scale, advancing the pursuit of generality, efficiency, and flexibility.
@@ 2 Related Work
Multimodal Embedding. These methods aim to project heterogeneous inputs, such as text and images, into a shared representation space for cross-modal understanding and retrieval (e.g Frome et al. (2013); Kiros et al. (2014); Faghri et al. (2018)). More recent large scale models such as CLIP (Radford et al., 2021), MetaCLIP (Xu et al., 2023; Chuang et al., 2025), BLIP (Li et al., 2022) and SigLIP (Zhai et al., 2023) encode each modality independently and apply contrastive training to enforce cross-modal alignment. More recent methods are built upon stronger VLMs (Xiao et al., 2024; Kong et al., 2025; Qin et al., 2025; Ju and Lee, 2025; Lin et al., 2025). For instance, VLM2Vec (Jiang et al., 2025) adapts Phi-3.5-V (Abdin et al., 2024), VLM2Vec-V2 (Meng et al., 2025) and GME (Zhang et al., 2024b) builds on Qwen2 (Wang et al., 2024) and LLaVE (Lan et al., 2025) finetunes on LLaVA (Li et al., 2024). Beyond architectural choices, the community has also explored innovative strategies in data construction and training. MegaPairs (Zhou et al., 2024) and mmE5 (Chen et al., 2025b) curate large-scale synthetic data with sophisticated pipelines to support contrastive training. UniME (Gu et al., 2025) achieves strong results through diverse data combined with teacher model distillation. B3 (Thirukovalluru et al., 2025b) incorporates novel insights into batch mining techniques. MoCa (Chen et al., 2025a) used continual pre-training to produce bidirectional embeddings. Nevertheless, many existing multimodal retrieval methods predominantly rely on single-vector retrieval, which hinders further scaling as embedding size becomes a bottleneck.
@@ 3.2 Our Design
Formally, let a VLM with parameters $\boldsymbol{\theta}$ define a conditional probability distribution $p_{\boldsymbol{\theta}}\left(\mathbf{y}\mid\mathbf{x},\mathcal{I}\right)$ where $\mathbf{x}=[x_{1},\dots,x_{n}]$ is the query or document text prompt and $\mathcal{I}$ are associated input images. MetaEmbed augments the input with learnable, input-specific Meta Tokens: queries use $\mathbf{M}_{q}\in\mathbb{R}^{R_{q}\times D}$, and candidates use $\mathbf{M}_{c}\in\mathbb{R}^{R_{c}\times D}$. For an input $(\mathbf{x},\mathcal{I})$, the transformer consumes $\mathbf{z}^{(0)}=\big[\;\mathbf{v}\ ;\ \mathbf{t}\ ;\ \mathbf{M}_{q}\ ;\ \mathbf{M}_{c}\;\big]\in\mathbb{R}^{(P+n+R_{q}+R_{c})\times D},$ where $\mathbf{v}$ and $\mathbf{t}$ are tokenized visual and text inputs. The model produces last-layer hidden states $\mathbf{H}=F_{\boldsymbol{\theta}}(\mathbf{z}^{(0)})\in\mathbb{R}^{(P+n+R_{q}+R_{c})\times D}$, where $F_{\boldsymbol{\theta}}$ denotes the transformer network parameterized by $\boldsymbol{\theta}$. We extract the final hidden states at the Meta Tokens positions to obtain query-side embeddings $\mathbf{E}_{\text{meta}}^{(q)}\in\mathbb{R}^{R_{q}\times D}$ or candidate-side embeddings $\mathbf{E}_{\text{meta}}^{(c)}\in\mathbb{R}^{R_{c}\times D}$, followed by L2 normalization. Each $\mathbf{E}_{\text{meta}}^{(q)}$ and $\mathbf{E}_{\text{meta}}^{(c)}$ constitutes a compact, contextualized multi-vector representation produced in two separate forward passes of the VLM.
@@ 4.2 Main Results
Table 2: NDCG@5 (%) results on the ViDoRe v2 benchmark, which covers 7 tasks on visual document retrieval. “Syn” denotes synthetic data, “Mul” indicates multilingual tasks, and “Bio” refers to biomedical domains. Bold denotes the best scores in the subset and the second-best scores are highlighted with underline.
@@ 4.3 Ablation Studies
How does the performance scales with the retrieval cost? We present the performance plots of different MetaEmbed models with respect to retrieval cost in Fig. 3. Data points in each plot correspond to MetaEmbed performance with a specific model size with varying test-time retrieval budgets. The dashed green line marks the best-performance single-vector retrieval model with identical training settings. The plots show that across model sizes, the curves rise steadily as more retrieval budget is allocated. While the improvement is modest for smaller models as MetaEmbed-3B shows +3.3 points relative gain against the single-vector method, we observe that it becomes more noticeable as the base model size grows and MetaEmbed brings the most pronounced improvements on the 32B model with a +6.6 points gain.

### v2

@@ Abstract
Universal multimodal embedding models have achieved great success in capturing semantic relevance between queries and candidates. However, current methods either condense queries and candidates into a single vector, potentially limiting the expressiveness for fine-grained information, or produce too many vectors that are prohibitive for multi-vector retrieval. In this work, we introduce MetaEmbed, a new framework for multimodal retrieval that rethinks how multimodal embeddings are constructed and interacted with at scale. During training, a fixed number of learnable Meta Tokens are appended to the input sequence. At test-time, their last-layer contextualized representations serve as compact yet expressive multi-vector embeddings. Through the proposed Matryoshka Multi-Vector Retrieval training, MetaEmbed learns to organize information by granularity across multiple vectors. As a result, we enable test-time scaling in multimodal retrieval where users can balance retrieval quality against efficiency demands by selecting the number of tokens used for indexing and retrieval interactions. Extensive evaluations on the Massive Multimodal Embedding Benchmark (MMEB) and the Visual Document Retrieval Benchmark (ViDoRe) confirm that MetaEmbed achieves state-of-the-art retrieval performance while scaling robustly to models with 32B parameters.
@@ 1 Introduction
Multimodal embedding models play an essential role in image search (Gordo et al., 2016), visual question answering (Hu et al., 2018; Zheng et al., 2021) and visual document retrieval (Faysse et al., 2025), where models project heterogeneous inputs into a unified representation space. While existing methods, including CLIP (Radford et al., 2021), BLIP (Li et al., 2022) and SigLIP (Zhai et al., 2023) have demonstrated superior performance in cross-modal retrieval, their performance remains limited in scenarios where the inputs involve complex and diverse instructions. Thanks to recent advances in building embeddings through foundation vision-language models (VLMs), one could apply contrastive learning on the extracted embedding from the hidden states of the last layer of a VLM to learn meaningful multimodal representations while retaining pre-trained knowledge.
@@
Despite growing progress in multimodal embedding VLMs, the common practice of condensing the entire query and candidate into a single vector is not an optimal choice, as fine-grained details are lost between modalities (Yao et al., 2022; Thrush et al., 2022) and this process has theoretical limitations (Weller et al., 2025). In text retrieval, ColBERT (Khattab and Zaharia, 2020) introduced a multi-vector late interaction mechanism that retains multiple token-level embeddings and uses a lightweight scoring between query and document token representations. This approach preserves significantly more contextual information than single-vector methods while still allowing independent encoding of queries and documents, and has motivated a recent trend of devising multi-vector embeddings for multimodal retrieval (Faysse et al., 2025; Xu et al., 2025; Günther et al., 2025).
However, multi-vector methods incur substantial efficiency costs in terms of index size, retrieval latency and feasibility. In these methods, each image is encoded into hundreds of patch embeddings, and each query text into several token embeddings, all of which must be stored and compared during retrieval. This results in large index sizes and slower retrieval processing. Moreover, multimodal-to-multimodal retrieval becomes impractical when both the query and candidate sides contain images, as similarity computation for each query-candidate pair involves interactions between thousands of query tokens and thousands of candidate tokens, making both training and inference prohibitive due to computational demands.
@@
Figure 1: Upper Left: Single vector retrieval method computes a score for each pair of query and candidate and uses a contrastive objective to maximize the score for corresponding pairs. Upper Right: Multi-vector retrieval aggregates maximum similarities across vector pairs before training. Lower: MetaEmbed structures query and candidate vectors into hierarchical nested groups and trains coarse-to-fine multi-vector embeddings that enable scalable and flexible retrieval.
@@
We first validate MetaEmbed on the Massive Multimodal Embedding Benchmark (MMEB) (Jiang et al., 2024) and Visual Document Retrieval Benchmarks (ViDoRe) v2 (Faysse et al., 2025; Macé et al., 2025), which represent a comprehensive suite of retrieval tests covering images, text and visual documents. Our experiments show that MetaEmbed achieves state-of-the-art retrieval performance across diverse scenarios. To further examine its generality and training scalability, we evaluate MetaEmbed with different VLM architectures and model sizes. Notably, test-time scalability remains effective even at 32B scale, with minimal diminishing returns as models grow larger. We hope MetaEmbed charts a path toward multimodal retrieval systems that are both accurate and deployable at scale, advancing the pursuit of generality, efficiency, and flexibility.
@@ 2 Related Work
Multimodal Embedding. These methods aim to project heterogeneous inputs into a shared representation space for cross-modal understanding and retrieval (e.g Frome et al. (2013); Kiros et al. (2014); Faghri et al. (2018)). More recent large scale models such as CLIP (Radford et al., 2021), MetaCLIP (Xu et al., 2023; Chuang et al., 2025), BLIP (Li et al., 2022) and SigLIP (Zhai et al., 2023) encode each modality independently and apply contrastive training to enforce cross-modal alignment. More recent methods are built upon stronger VLMs (Xiao et al., 2024; Kong et al., 2025; Qin et al., 2025; Ju and Lee, 2025; Lin et al., 2025; xiao2025locore). For instance, VLM2Vec (Jiang et al., 2025) adapts Phi-3.5-V (Abdin et al., 2024), VLM2Vec-V2 (Meng et al., 2025) and GME (Zhang et al., 2024b) builds on Qwen2 (Wang et al., 2024) and LLaVE (Lan et al., 2025) finetunes on LLaVA (Li et al., 2024). Beyond architectural choices, the community has also explored innovative strategies in data construction and training. MegaPairs (Zhou et al., 2024) and mmE5 (Chen et al., 2025b) curate large-scale synthetic data with sophisticated pipelines to support contrastive training. UniME (Gu et al., 2025) achieves strong results through diverse data combined with teacher model distillation. B3 (Thirukovalluru et al., 2025) incorporates novel insights into batch mining techniques. MoCa (Chen et al., 2025a) used continual pre-training to produce bidirectional embeddings. Nevertheless, many existing multimodal retrieval methods predominantly rely on single-vector retrieval, which hinders further scaling as embedding size becomes a bottleneck.
@@ 3.2 Our Design
Formally, let a VLM with parameters $\boldsymbol{\theta}$ define a conditional probability distribution $p_{\boldsymbol{\theta}}\left(\mathbf{y}\mid\mathbf{x},\mathcal{I}\right)$ where $\mathbf{x}=[x_{1},\dots,x_{n}]$ is the query or document text prompt and $\mathcal{I}$ are associated input images. MetaEmbed augments the input with learnable Meta Tokens: queries use $\mathbf{M}_{q}\in\mathbb{R}^{R_{q}\times D}$, and candidates use $\mathbf{M}_{c}\in\mathbb{R}^{R_{c}\times D}$. For an input $(\mathbf{x},\mathcal{I})$, the transformer consumes $\mathbf{z}^{(0)}=\big[\;\mathbf{v}\ ;\ \mathbf{t}\ ;\ \mathbf{M}_{q}\ ;\ \mathbf{M}_{c}\;\big]\in\mathbb{R}^{(P+n+R_{q}+R_{c})\times D},$ where $\mathbf{v}$ and $\mathbf{t}$ are $P$ visual patches tokens and text inputs. The model produces last-layer hidden states $\mathbf{H}=F_{\boldsymbol{\theta}}(\mathbf{z}^{(0)})\in\mathbb{R}^{(P+n+R_{q}+R_{c})\times D}$, where $F_{\boldsymbol{\theta}}$ denotes the transformer network parameterized by $\boldsymbol{\theta}$. We extract the final hidden states at the Meta Tokens positions to obtain query-side embeddings $\mathbf{E}_{\text{meta}}^{(q)}\in\mathbb{R}^{R_{q}\times D}$ or candidate-side embeddings $\mathbf{E}_{\text{meta}}^{(c)}\in\mathbb{R}^{R_{c}\times D}$, followed by L2 normalization. Each $\mathbf{E}_{\text{meta}}^{(q)}$ and $\mathbf{E}_{\text{meta}}^{(c)}$ constitutes a compact, contextualized multi-vector representation produced in two separate forward passes of the VLM.
@@ 4.2 Main Results
Table 2: NDCG@5 (%) results on the ViDoRe v2 benchmark, which covers 7 tasks on visual document retrieval. “Syn” denotes synthetic data, “Mul” indicates multilingual tasks, and “Bio” refers to biomedical domains.
@@ 4.3 Ablation Studies
How does the performance scale with the retrieval cost? We present the performance plots of different MetaEmbed models with respect to retrieval cost in Fig. 3. Data points in each plot correspond to MetaEmbed performance with a specific model size with varying test-time retrieval budgets. The dashed green line marks the best-performance single-vector retrieval model with identical training settings. The plots show that across model sizes, the curves rise steadily as more retrieval budget is allocated. While the improvement is modest for smaller models as MetaEmbed-3B shows +3.3 points relative gain against the single-vector method, we observe that it becomes more noticeable as the base model size grows and MetaEmbed brings the most pronounced improvements on the 32B model with a +6.6 points gain.